Write a system of equations for the given information, and then use the system to find the soultion or solutions..
a math teacher wants to buy 25 new scientific calculators for the classroom. casio and zenith make two different types of calculators. casio calculators cost $15.00 each. zenith calculators cost $12.00 each. the math teacher has $330.00 to spend. how many of each calculator can she buy?
THANKS FOR HELP!!
Lynn

Answer by ReadingBoosters(3246)

15c + 12z = 330
c+z=25 : rearrange as c = 25 - z
...
15(25-z) + 12z = 330
375 - 15z + 12z = 330
-3z = 330 - 375
-3z = -45
z=15 zenith calculators
c = 25-z = 25-15 = 10 casio calculators
.....................
Have more questions? HomeworkHelpers@readingboosters.com
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.